OpenStudy (anonymous):

Simplify. x − 3y − (−2x + 5y)

OpenStudy (anonymous):

3x-8y

OpenStudy (anonymous):

SHOW WORK

OpenStudy (anonymous):

\[x − 3y − (−2x + 5y)\] \[=x-3y+2x-5y\] \[=3x-8y\]